OpenStudy (anonymous):

help with absolute value questions. Ill give medal and become fan if u help with all! Its 6 questions for 8th grade

OpenStudy (anonymous):

Solve |2y - 2| = y.

OpenStudy (anonymous):

you have to split them into two equations then solve. For |2n-3|=5 you write 2n-3 = 5 and -(2n-3) = 5 then solve for n. so 2n-3 = 5 then 2n = 5 + 3, 2n = 8, then n=4 now we solve for -(2n-3) =5 so we have -2n + 3 = 5 then -2n = 2 then n = -1 so our solution is n= 4 and n = -1

OpenStudy (anonymous):

do u think u can help with a few more? just two more? im really desperate!

OpenStudy (anonymous):

ok, but did you understand my explanation of the first one? Try the other ones one more time following my example of the first one. I'll help you if you get stuck.

OpenStudy (anonymous):

ok so im doing the 3u one for the postive one i got 6 ill solve the negative side now

OpenStudy (anonymous):

im stuck on the negative

OpenStudy (anonymous):

where specifically are you getting stuck?

OpenStudy (anonymous):

well isnt the answer -6

OpenStudy (anonymous):

and 6

OpenStudy (anonymous):

the answer is 6 for both of them

OpenStudy (anonymous):

posisitve 6 or negative

OpenStudy (anonymous):

positive 6 for both

OpenStudy (anonymous):

ok Last one! plz!

OpenStudy (anonymous):

Solve |2y - 2| = y.

OpenStudy (anonymous):

idk where to even start

OpenStudy (anonymous):

@ktm

OpenStudy (anonymous):

ill help

OpenStudy (anonymous):

2y-2 = y then 2y-y = 2 so then we have y = 2

OpenStudy (anonymous):

2 and what other number

OpenStudy (anonymous):

this is easy, im in 8th grade too :)

OpenStudy (anonymous):

then for the negative we have -(2y-2)=y so -2y + 2 = y then 2 = 3y so y = 2/3

OpenStudy (anonymous):

-2y-2=y -2=3y -2/3=y

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

which school do u got to

OpenStudy (anonymous):

im sorry i am wrong,didn't take count of the absolute value, i go to connections academy

OpenStudy (anonymous):

omg i go to online school too! laurel springs tho... i play tennis so i need flxible schedule

OpenStudy (anonymous):

i think alot of people in OS go to online school.

OpenStudy (anonymous):

yup, don't forget to multiply all the terms in the brackets by the negative sign -(2y-2) = y becomes -2y+2 = y